# MOKA: Open-Vocabulary Robotic Manipulation through Mark-Based Visual Prompting

### v1

@@ B-D Motion Execution
We can decompose a manipulation subtask into an optional grasping phase and a manipulation phase. For some tasks like tool using, the robot needs to first grasp an object before making contact with another object. For other tasks where the robot can directly interact with target objects (e.g. pushing, pressing button), the “grasp__\__keypoint” field will be empty. For such tasks, we will skip the execution of grasping phase.
Grasping phase. We first sample 30 antipodal 4DoF grasp proposals based on the primary camera’s depth image, which is cropped by the bounding box of object_normal-_\__grasped. Here, we use the same antipodal grasp proposal method used in DexNet 2.0 [29]. We then lift the VLM-predicted grasp point from 2D image to the robot frame based on the primary camera’s intrinsic and extrinsic parameters. After that, we can select the nearest grasp proposal based on the lifted grasp point, and execute the 4DoF grasp on the robot.
Manipulation phase. After obtaining the pre-contact tile and post-contact tile, we first sample pre-contact waypoint and post-contact waypoint from the predicted tiles. Since the waypoints are in the free space, its 3D position cannot be determined given its 2D projection on the image. So we also assign the VLM-predicted height to them. We only consider the cases where the waypoints are in same height with target point (e.g. pushing), or above the target point (e.g. placing) in most common table top manipulation scenarios.
@@ B-E Policy Distillation
We employ the base checkpoint of Octo [35] and perform full-model finetuning using their official instruction at https://github.com/octo-models/octo. The default architecture of Octo is a transformer-based diffusion policy, which is conditioned on a task discription or goal image through a task tokenizer along with tokenized pixel observations from multiple cameras, and predicts actions through a diffusion head. The diffusion head consists of a 3-layer MLP with a hidden dimension of 256 using a standard DDPM objective [13].
We mostly reuse the hyperparameters as in Octo [35]. For customized hyper-parameters we used, please refer to Tab. VIII.
@@ B-F Evaluation
We evaluate MOKA in both zero-shot and in-context learning settings. For the zero-shot setting, we keep all the above prompts unchanged but only change the language task description (e.g. “sweep the garbage”, “pick up the perfume”, “press the button”).
@@ C-A Ablation Study
Figure 8: Qualitative results of ablation studies. Without keypoint description or chain-of-thought prompting, the model can make mistakes in keypoint prediction (left column) and waypoint prediction (right column).
C-B Additional Tasks
For our four additional tasks, please refer to our supplementary video.

### v3

@@ B-D Motion Execution
We can decompose a manipulation subtask into an optional grasping phase and a manipulation phase. For some tasks like tool using, the robot needs to first grasp an object before making contact with another object. For other tasks where the robot can directly interact with target objects (e.g., pushing, pressing button), the “grasp__\__keypoint” field will be empty. For such tasks, we will skip the execution of grasping phase.
Grasping phase. We first sample 30 antipodal 4-DoF grasp proposals based on the primary camera’s depth image, which is cropped by the bounding box of object__\__grasped. Here, we use the same antipodal grasp proposal method used in DexNet 2.0 [39]. We then lift the VLM-predicted grasp point from 2D image to the robot frame based on the primary camera’s intrinsic and extrinsic parameters. After that, we can select the nearest grasp proposal based on the lifted grasp point, and execute the 4-DoF grasp on the robot.
Manipulation phase. After obtaining the pre-contact tile and post-contact tile, we first sample pre-contact waypoint and post-contact waypoint from the predicted tiles. Since the waypoints are in the free space, its 3D position cannot be determined given its 2D projection on the image. So we also assign the VLM-predicted height to them. We only consider the cases where the waypoints are in same height with target point (e.g., pushing), or above the target point (e.g., placing) in most common table top manipulation scenarios.
@@ B-E Policy Distillation
We employ the base checkpoint of Octo [46] and perform full-model finetuning using their official instruction. The default architecture of Octo is a transformer-based diffusion policy, which is conditioned on a task discription or goal image through a task tokenizer along with tokenized pixel observations from multiple cameras, and predicts actions through a diffusion head. The diffusion head consists of a 3-layer MLP with a hidden dimension of 256 using a standard DDPM objective [19].
We mostly reuse the hyperparameters as in Octo [46]. For customized hyper-parameters we used, please refer to TABLE VIII.
@@ B-F Evaluation
We evaluate MOKA in both zero-shot and in-context learning settings. For the zero-shot setting, we keep all the above prompts unchanged but only change the language task description (e.g., “sweep the garbage”, “pick up the perfume”, “press the button”).
@@ C-A Ablation Study
Figure 8: Qualitative results of ablation studies. Without the keypoint description or chain-of-thought prompting, the model can make mistakes in keypoint prediction (left column) and waypoint prediction (right column).
C-B Additional Robustness Analysis
We provide additional results for robustness analysis. The predicted affordances for the drawer closing and fur removal tasks with various objects and initial poses are shown in Fig. 10 and Fig. 10. In these two figures, each column in the image uses similar initial arrangements of objects, while the two rows use different objects. For the drawer closing tasks, MOKA can consistently predict the waypoints and target points following the correct closing motion regardless of the poses and appearances of drawers and other distractors. For the fur removal tasks, MOKA can predict all the keypoints correctly with a feasible motion in different environments.
Figure 9: Robustness analysis for the drawer closing task.
Figure 10: Robustness analysis for the fur removal task.